Extra Credit 56. Who are the payers for healthcare in the U.S.? A. Medicare. B. Medicaid. Private health insurance. All of the above. 57. The Affordable Care Act (Obamacare) has fixed the following problems EXCEPT: A. the uninsured problem as more people gain health insurance coverage. B. discrimination problem as people with pre-existing conditions allowed to sign up for health insurance. D, the adverse selection problem due to the individual mandate. ng cost problem. II. Short Answer and Graph Questions- write your answers neatly and clearly. 1. The marginal cost is MC- 10 and total cost is TC 10Q. The demand curve is Q 20-P. Given the information, answer the following questions. Show your work to receive full credit (14 points). a. For a perfectly competitive firm: i) Calculate the equilibrium quantity: 20-2 ii) Calculate the equilibrium price: ii) Calculate the maximizing profit:
Answer #1

56) Answer: All the above

The payers for health care in US is a combination of both private and public health insurance. Medicare and Medicaid comes under public health insurance.

1)

TC = 10Q

MC = 10

Q = 20-P

Where P = 20-Q

Total revenue = TR = P*Q = (20-Q)*Q

TR = 20Q-Q2

Marginal revenue = MR = d(TR)/dQ

MR = d(20Q-Q2)/dQ

MR = 20-2Q

To find equilibrium price and quantity

MR = MC

20-2Q = 10

2Q = 10

Q = 5 units

P = 20-Q = 20-5 =15

TC = 10*5 =50

TR = 15*5 =75

Maximizing profit = 75-50 = 25